# KinDEL: DNA-Encoded Library Dataset for Kinase Inhibitors

### v1

@@ 5.2 Computational Methods on DEL Data
Computational methodology for DEL data is still a nascent area, largely due to a lack of publicly available data. Current works typically directly apply fingerprint and graph neural network approaches that are popular within property prediction models. Due to the noisiness intrinsic to the data, many methods bin the labeled data into a binary classification to avoid overfitting on the raw data. For instance, McCloskey et al. (2020) trained classification models on disynthons using variants of typical graph neural networks. This approach was applied by Ahmad et al. (2023) to find first-in-class WD Repeat Domain 91 ligands, using molecules from buyable catalogues. Torng et al. (2023) follows up on this work and uses Weave GCN for hit identification of CA-IX. However, the true binding process is observed on a discrete scale, so more recent works have focused on building probabilistic models of binding. Lim et al. (2022) proposed a new uncertainty-based loss function for training regression models, and demonstrated the efficacy of their method using various graph neural network models. Shmilovich et al. (2023) takes this process one step further and incorporates 3-D docked poses to leverage the scale of DEL data by using multiple-instance learning to learn over poses. Other works exploit compositionality of DELs as inductive biases of the models. Binder et al. (2022) tries to explicitly model the partial products, while Chen et al. (2024) computes a hierarchical representation of DEL molecules and incorporates a probabilistic loss. Koziarski et al. (2024) explores generative models of DEL molecules using GFlowNets. Many of these developments are relatively new, and we hope that KinDEL will enable the further development of these methodologies in the field.
@@ 6 Conclusion
DELs have emerged as a high-throughput technology that enables screenings of large combinatorial small molecule libraries. However, DEL data has many sources of intrinsic noise stemming from synthesis and selection experiments, necessitating the right machine learning tools to extract the correct signal in the data. We introduce KinDEL as a 81M molecule dataset with selection data from two targets, MAPK14 and DDR1, in order to highlight the ability to model DEL data to find potent binders. Our DEL is built with chemical diversity in mind, with molecules in the drug-like domain. We additionally release biophysical data for molecules synthesized both on- and off-DNA to validate our models trained on DEL data. We hope that our public data release and benchmarking will engender more interest in DEL data as an important chemical modality for machine learning research in the future.
7 Reproducibility Statement
We have included both experimental details of the data generation as well as all the model details for our benchmarking tasks in Appendix A and D. The code to replicate our experiments and data can be found at: https://github.com/insitro/kindel. See the README file in the code for more details.
@@ A.2 DEL Selection
From the 10% of eluates conserved from round 2 and the round 3 eluate were taken 5 µL of sample that was subsequently PCR amplified in 50 µL reactions using New England BioLabs 2x Q5 mastermix (Cat. #M0492S) and custom designed barcode primers to differentiate both the round, target and experiment. Following this 1 µL of the amplified material is transferred to a 25 µL PCR reaction with Q5 and custom P5 and P7 primers to install the relevant sequencing primers. The PCR mixture is purified using a QIAquick PCR Purification Kit (Cat. #28104), this sample is then further purified to remove any DNA products of incorrect size using a PippenHT. The purified samples are submitted to next-generation sequencing utilizing the Novaseq platform with a 2x150 BP S4 kit.
@@ A.3 Biophysical Assays
hDDR1-1 (593-913, 730-735) gene was cloned into pFastBac1 expression vector by PCR. An N-terminal His-TEV tag was added by using the forward primer 5’-GGTCCGAAGCGCGCGGAATTCACCATGCACCATCACCATCACCAC-3’ and the reverse primer 5’-ACCTTGGAAGTACAGGTTCTC-3’. The forward primer included the restriction site EcoRI. The hMAPK14 (593-913, 730-735) gene was cloned using the forward primer 5’-GAGAACCTGTACTTCCAAGGTCCTGGTGCTGTGGGTGACGGT-3’ and the reverse primer 5’-GGCTCTAGATTCGAAAGCGGCCGCTTACACAGTGTTCAGAGCGTC-3’. The reverse primer included the restriction site NotI. This generated plasmid pFastBac1-His-TEV-hDDR1-1 [593-913 (730-735), WT].
hDDR1-2 (593-913, 730-735) gene was cloned into pFastBac1 expression vector by PCR. An N-terminal His-TEV-Avi tag was added by using the forward primer 5’-GGTCCGAAGCGCGCGGAATTCACCATGCATCATCACCATCACCAC-3’ and the reverse primer 5’-TTCGTGCCATTCAATTTTCTG-3’. The forward primer included the restriction site EcoRI. The hMAPK14 (593-913, 730-735) gene was cloned using the forward primer 5’-CTCAGAAAATTGAATGGCACGAACCTGGTGCTGTGGGTGACGGT-3’ and the reverse primer 5’-GGCTCTAGATTCGAAAGCGGCCGCTTACACAGTGTTCAGAGCGTC-3’. The reverse primer included the restriction site NotI. This generated plasmidpFastBac1-His-TEV-Avi-hDDR1-2 [593-913 (730-735), WT].
All plasmids were amplified using DH5 cells followed by DNA extraction. Insertion of the genes were verified by sequencing.
Protein Production
@@
Assay setup. In a 384 well black small volume microplate (Greiner Bio-One, 784900), the annealed compounds were dispensed using Echo650 (Labcyte) at a constant final concentration of 4 nM in 137 mM NaCl, 2.7 mM KCl, 9.8 mM Phosphate buffer, 0.01% Tween-20. MAPK14-2 and DDR1-1 were serially diluted 1:1 starting from a top dose of in 137 mM NaCl, 2.7 mM KCl, 9.8 mM Phosphate buffer, 0.01% Tween-20 for a total 16-point dilution and were transferred to the 384 well black small volume microplate containing the annealed compounds using Agilent Bravo G5563A. The final assay concentration was 2 nM the annealed compound and top dose of the protein. The final assay volume was . The serially diluted protein was incubated with the annealed DEL compounds and then the fluorescence polarization was measured using PerkinElmer EnVision 2105 in milli-P (mP), where mP = 1000(S-GP)/(S+GP) (S and P are background subtracted fluorescence count rates, and G is an instrument and assay dependent factor). The dose-response curves (mP vs. [Protein]) were fit in GraphPad Prism using the equation, Y = Bmax X / (KD+X) +NSX+Background, where Y is mP, X is [Protein], Bmax is the maximum specific binding in the same units as Y, KD is the equilibrium dissociation constant, in the same units as X, NS is the slope of nonspecific binding in Y units divided by X units, Background is the amount of nonspecific binding with no added radioligand.
@@
Compounds were tested using Surface Plasmon Resonance (SPR) using a Biacore T200 and a Biacore S200 (Cytiva Life Sciences). N-terminally Avi-tagged DDR1 [593-913 (730-735)] and N-terminally Avi-tagged MAPK14 (1-360) were biotinylated enzymatically and immobilized to a streptavidin coated SPR chip (Series S SA, Cytiva, 29699621). Approximately 4400 RU of hDDR1-2 and 4900 RU of MAPK14 were immobilized for dose response assays using 1xHBS-P+ (10 mM HEPES, 150 mM NaCl, 0.05% v/v Surfactant P20; pH 7.4). 1xHBS-P+ was prepared by mixing 10xHBS-P+ (Cytiva, BR100671) with HPLC-grade water (Fisher Scientific, W51).
Compounds were diluted in the SPR running buffer which consisted of 1xHBS-P+ supplemented with 5% DMSO (VWR, 76177-938). Multi-cycle kinetics was used to determine compound affinities. Compounds were injected in a series of increasing concentrations with a /min flow rate, a 60 s contact time and a 300 s dissociation time. Compound sensorgrams were processed and fit using Biacore Evaluation software, with DMSO correction applied during sensorgram processing. Binding and dissociation curves for each compound were globally fit to obtain an on- and off-rate kinetic constant (ka and kd, respectively) which were used to determine the overall binding constant (KD) for each compound.
@@ Appendix B Experimental Replicability
DEL experiments suffer from the substantial noise that stems from the large scale of combinatorial libraries screened simultaneously within a single tube. Key sources of this noise include inconsistent control over the quantities of each compound in the mixture, synthesis-related challenges such as the introduction of side products and low synthetic yields, as well as errors during DNA sequencing. Figure 6 illustrates the strong correlation between experimental replicates in our DEL experiments, underscoring the reproducibility of the panning procedures we conducted.
@@ Appendix C Validation Set Selection
We selected validation compounds to cover a range of values and chemical diversity, allowing us to better assess the models’ ability to rank molecules across diverse binding affinities. To achieve this, we used an ensemble of models described in Section 3, and then clustered the molecules based on chemical similarity via Butina clustering using Morgan fingerprints. This constituted the molecules in the “in-library” set. We also surveyed literature for tool compounds, which we also re-synthesized on-DNA. This made up the additional molecules found in the “extended” set.

### v2

@@ 5.1 Current datasets
The AIRCHECK database facilitates the public release of large-scale DEL datasets, notably including a 3-billion-molecule library screened against WDR91 (Wellnitz et al., 2024). However, chemical structures are masked with fingerprints in these commercial-sized libraries, which inherently limits the range of applicable modeling approaches.
@@ 5.2 Computational Methods on DEL Data
Computational methodology for DEL data is still a nascent area, largely due to a lack of publicly available data. Current works typically directly apply fingerprint and graph neural network approaches that are popular within property prediction models. Due to the noisiness intrinsic to the data, many methods bin the labeled data into a binary classification to avoid overfitting on the raw data. For instance, McCloskey et al. (2020) trained classification models on disynthons using variants of typical graph neural networks. This approach was applied by Ahmad et al. (2023) to find first-in-class WD Repeat Domain 91 ligands, using molecules from buyable catalogues. Torng et al. (2023) follows up on this work and uses Weave GCN for hit identification of CA-IX. However, the true binding process is observed on a discrete scale, so more recent works have focused on building probabilistic models of binding. Lim et al. (2022) proposed a new uncertainty-based loss function for training regression models, and demonstrated the efficacy of their method using various graph neural network models. (Shmilovich et al., 2023) takes this process one step further and incorporates 3-D docked poses to leverage the scale of DEL data by using multiple-instance learning to learn over poses. Other works exploit compositionality of DELs as inductive biases of the models. (Binder et al., 2022) tries to explicitly model the partial products, while (Chen et al., 2024) computes a hierarchical representation of DEL molecules and incorporates a probabilistic loss. (Koziarski et al., 2024) explores generative models of DEL molecules using GFlowNets. Many of these developments are relatively new, and we hope that KinDEL will enable the further development of these methodologies in the field.
@@ 6 Conclusion
DELs have emerged as a high-throughput technology that enables screenings of large combinatorial small molecule libraries. However, DEL data has many sources of intrinsic noise stemming from synthesis and selection experiments, necessitating the right machine learning tools to extract the correct signal in the data. We introduce KinDEL as a 81M molecule dataset with selection data from two targets, MAPK14 and DDR1, in order to highlight the ability to model DEL data to find potent binders. Our DEL is built with chemical diversity in mind, and many of the molecules in the library have properties within the range of approved drugs. We additionally release biophysical data for molecules synthesized both on- and off-DNA to validate our models trained on DEL data. We hope that our public data release and benchmarking will engender more interest in DEL data as an important chemical modality for machine learning research in the future.
Impact Statement
The broader impact of this work includes the potential to accelerate the identification of novel therapies, ultimately benefiting public health. However, as with many datasets in the chemical and biological sciences, there is a potential risk of misuse, such as designing harmful substances. While there are inherent societal implications with advancing drug discovery, we do not anticipate any specific ethical concerns that must be highlighted beyond the established consequences of advancing machine learning techniques in this domain.
@@ A.2 DEL Selection
From the 10% of eluates conserved from round 2 and the round 3 eluate were taken 5 µL of sample that was subsequently PCR amplified in 50 µL reactions using New England BioLabs 2x Q5 mastermix (Cat. #M0492S) and custom designed barcode primers to differentiate both the round, target and experiment. Following this 1 µL of the amplified material is transferred to a 25 µL PCR reaction with Q5 and custom P5 and P7 primers to install the relevant sequencing primers. The PCR mixture is purified using a QIAquick PCR Purification Kit (Cat. #28104), this sample is then further purified to remove any DNA products of incorrect size using a PippenHT. The purified samples are submitted to next-generation sequencing utilizing the Novaseq platform with a 2x150 BP S4 kit. The sequencing depth (number of read counts/sequence counts) is 514.5M/65.8M for MAPK14, 296.5M/49.8M for DDR1, and 440.7M/52M for BCA in all three replicates.
@@ A.3.1 Molecular Biology
hDDR1-1 (593-913, 730-735) gene was cloned into pFastBac1 expression vector by PCR. An N-terminal His-TEV tag was added by using the forward primer 5’-GGTCCGAAGCGCGCGGAATTCACCATGCACCATCACCATCACCAC-3’ and the reverse primer 5’-ACCTTGGAAGTACAGGTTCTC-3’. The forward primer included the restriction site EcoRI. The hDDR1-1 (593-913, 730-735) gene was cloned using the forward primer 5’-GAGAACCTGTACTTCCAAGGTCCTGGTGCTGTGGGTGACGGT-3’ and the reverse primer 5’-GGCTCTAGATTCGAAAGCGGCCGCTTACACAGTGTTCAGAGCGTC-3’. The reverse primer included the restriction site NotI. This generated plasmid pFastBac1-His-TEV-hDDR1-1 [593-913 (730-735), WT].
hDDR1-2 (593-913, 730-735) gene was cloned into pFastBac1 expression vector by PCR. An N-terminal His-TEV-Avi tag was added by using the forward primer 5’-GGTCCGAAGCGCGCGGAATTCACCATGCATCATCACCATCACCAC-3’ and the reverse primer 5’-TTCGTGCCATTCAATTTTCTG-3’. The forward primer included the restriction site EcoRI. The hDDR1-2 (593-913, 730-735) gene was cloned using the forward primer 5’-CTCAGAAAATTGAATGGCACGAACCTGGTGCTGTGGGTGACGGT-3’ and the reverse primer 5’-GGCTCTAGATTCGAAAGCGGCCGCTTACACAGTGTTCAGAGCGTC-3’. The reverse primer included the restriction site NotI. This generated plasmidpFastBac1-His-TEV-Avi-hDDR1-2 [593-913 (730-735), WT].
BCA2-1 (1-260, WT) gene was cloned into pET28a expression vector by PCR. An N-terminal His-TEV-Avi tag was added by using the forward primer 5’-CTTTAAGAAGGAGATATACCATGGGCCATCATCACCATCACCAC-3’ and the reverse primer 5’-TTCGTGCCATTCAATTTTCTG-3’. The forward primer included the restriction site NcoI. The BCA2-1 (1-260, WT) gene was cloned using the forward primer 5’-CTCAGAAAATTGAATGGCACGAAATGAGCCATCATTGGGGCTATGGCAAAC-3’ and the reverse primer 5’-GTGGTGGTGGTGGTGGTGCTCGAGTTATTTCGGAAAGCCGCGCACT-3’. The reverse primer included the restriction site XhoI. This generated plasmid pET28a-His-TEV-Avi-BCA2(1-260,WT).
All cloning was performed using the Beyotime Seamless Cloning kit (D7010M). All plasmids were amplified using DH5 cells followed by DNA extraction. Insertion of the genes were verified by sequencing.
A.3.2 Protein Production
MAPK14
@@ A.3.2 Protein Production
BCA2
The BCA plasmid, pET28a-His-TEV-Avi-BCA2(1-260,WT), was transformed into BL21-Gold (DE3) competent cells (Agilent, 230132) and plated on LB/agar/kanamycin (50 µg/mL) medium then left to grow at 37°C. Fresh colonies from transformed BL21-Gold cells were picked and used to inoculate 100 mL of LB medium (10 g tryptone, 10 g NaCl, 5 g yeast extract per liter water) supplemented with 50 µg/mL kanamycin and cultured overnight at 220 rpm 37°C. The overnight culture was added to 10 L LB/kanamycin (50 µg/mL) and grown at 180 rpm 37°C until OD600 reached 0.600. Isopropyl ß-D-1-thiogalactopyranoside (IPTG) was then added (final concentration 0.3 mM) and the culture was left to continue to grow at 16°C, 160 rpm overnight.
Cells were harvested the following day at 8,000 rpm at 4°C and stored at -20°C. Cell pellets were resuspended in lysis buffer (50 mM Tris, 500 mM NaCl, 1 mM TCEP, 1 mM PMSF, 10% glycerol, pH 8.0 and 100 U/mL Benzonase) and lysed using a high pressure cell homogenizer. Lysate was centrifuged at 13,000 rpm for 30 minutes at 4°C, twice, to remove cell debris. Lysate supernatant was then applied to 5 mL of Ni-Resin (which had been pre-equilibrated with lysis buffer) and the lysate/Ni mixture was incubated for 1 hr at 4°C. Resin was then loaded into a column and the column was washed with wash buffer (50 mM Tris, 500 mM NaCl, 1 mM TCEP, 10% glycerol, pH 8.0) and eluted with elution buffer (50 mM Tris, 500 mM NaCl, 1 mM TCEP, 10% glycerol, pH 8.0 and 250 mM imidazole). Eluted protein was cleaved with TEV protease at a 1:10 (w/w) ratio and dialyzed against dialysis buffer (50 mM Tris, 100 mM NaCl, 1 mM TCEP, 7.5 mM MgCl2, 5% glycerol, pH 8.0) at 4°C overnight.
The dialyzed protein was incubated at 18°C for 4 hrs with ATP (1 mM), Biotin (0.5 mM), and BirA (200 nM), prior to loading onto Ni-resin to remove cleaved hexahistidine tags. Protein was loaded onto a Ni-resin column, pre-equilibrated with lysis buffer and the cleaved protein was collected in the flow-through (cleaved-tag left on the column).
Protein was collected and concentrated with a 30 kDa Millipore Amicon Ultra-15 centrifugal filter unit. Concentrated protein was loaded onto a Superdex 75 column pre-equilibrated with SEC buffer. Protein was collected according to UV-vis signal and follow-up SDS-PAGE verification. Protein fractions were pooled and concentrated with a 30 kDa Millipore Amicon Ultra-15 centrifugal filter unit and exchanged into a final buffer containing 25 mM HEPES, 200 mM NaCl, 1 mM TCEP at pH 7.5. A final purified protein yield of 8.38 mg/L cell culture was obtained.
BCA used for fluorescence polarization assay was obtained from Sigma-Aldrich (Product No. C2522).
A.3.3 Biophysical Methods
@@
Assay setup. In a 384 well black small volume microplate (Greiner Bio-One, 784900), the annealed compounds were dispensed using Echo650 (Labcyte) at a constant final concentration of 4 nM in 137 mM NaCl, 2.7 mM KCl, 9.8 mM Phosphate buffer, 0.01% Tween-20. MAPK14-2, DDR1-1, and BCA were each serially diluted 1:1 starting from a top dose of in 137 mM NaCl, 2.7 mM KCl, 9.8 mM Phosphate buffer, 0.01% Tween-20 for a total 16-point dilution and were transferred to the 384 well black small volume microplate containing the annealed compounds using Agilent Bravo G5563A. The final assay concentration was 2 nM the annealed compound and top dose of the protein. The final assay volume was . The serially diluted protein was incubated with the annealed DEL compounds and then the fluorescence polarization was measured using PerkinElmer EnVision 2105 in milli-P (mP), where ( and are background subtracted fluorescence count rates, and is an instrument and assay dependent factor). The dose-response curves ( vs. [Protein]) were fit in GraphPad Prism using the equation, , where is , is [Protein], is the maximum specific binding in the same units as , is the equilibrium dissociation constant, in the same units as , is the slope of nonspecific binding in units divided by units, is the amount of nonspecific binding with no added radioligand.
@@
Compounds were tested using Surface Plasmon Resonance (SPR) using a Biacore T200 and a Biacore S200 (Cytiva Life Sciences). N-terminally biotinylated DDR1 [593-913 (730-735)], MAPK14 (1-360), and BCA2-1 (1-260, WT) were immobilized to a streptavidin coated SPR chip (Series S SA, Cytiva, 29699621). Approximately 4400 RU of hDDR1-2, 4900 RU of MAPK14-1, and 3300 RU of BCA2-1 were immobilized for dose response assays using 1xHBS-P+ (10 mM HEPES, 150 mM NaCl, 0.05% v/v Surfactant P20; pH 7.4). 1xHBS-P+ was prepared by mixing 10xHBS-P+ (Cytiva, BR100671) with HPLC-grade water (Fisher Scientific, W51).
Compounds were diluted in the SPR running buffer which consisted of 1xHBS-P+ supplemented with 5% DMSO (VWR, 76177-938). Multi-cycle kinetics was used to determine compound affinities. Compounds were injected in a series of increasing concentrations with a /min flow rate, a 60 s contact time and a 300 s dissociation time. Compound sensorgrams were processed and fit using Biacore Evaluation software, with DMSO correction applied during sensorgram processing. Binding and dissociation curves for each compound were globally fit to obtain an on- and off-rate kinetic constant (k and k, respectively) which were used to determine the overall binding constant (K) for each compound.
@@ Appendix B Experimental Replicability
DEL experiments suffer from the substantial noise that stems from the large scale of combinatorial libraries screened simultaneously within a single tube. Key sources of this noise include inconsistent control over the quantities of each compound in the mixture, synthesis-related challenges such as the introduction of side products and low synthetic yields, as well as errors during DNA sequencing. Figure LABEL:fig:correlation-plots illustrates the strong correlation between experimental replicates in our DEL experiments, underscoring the reproducibility of the panning procedures we conducted.
@@ Appendix C Validation Set Selection
We selected validation compounds to cover a range of values and chemical diversity, allowing us to better assess the models’ ability to rank molecules across diverse binding affinities. To achieve this, we used an ensemble of models described in Section 3, and then clustered the molecules based on chemical similarity via Butina clustering using Morgan fingerprints. This constituted the molecules in the “in-library” set. Figure LABEL:fig:kd-dist-heldout shows the distributions of the pKD values of the selected compounds. We also surveyed literature for tool compounds, which we also re-synthesized on-DNA. This made up the additional molecules found in the “extended” set.
@@ Appendix D Model Hyperparameters
Chemprop uses three layers of bond message passing with the hidden dimension of 300. Then, the information is aggregated using a global average pooling layer, and one linear layer is used to make predictions. The model is trained using the default Chemprop optimization procedure, which employs the Adam optimizer with a Noam learning rate scheduler and 2 epochs of warmup.
@@
Appendix E Dataset Splitting
It is often difficult ensure that information from the test set does not leak into train. One popular method is calculating Bemis-Murcko scaffolds (Bemis & Murcko, 1996) for each compound in a dataset and assigning compounds to a split by scaffold. Unfortunately, this method often fails for DEL data. Below we calculate Bemis-Murcko scaffolds for the top million compounds in the MAPK14 dataset using rdkit (Landrum, 2010). There are 300,000 unique scaffolds calculated for these compounds, implying many compounds have a unique scaffold (see Fig. LABEL:fig:murckoa). Additionally the common scaffolds are often trivial (see Fig. LABEL:fig:murckob). The most common scaffold is a simple benzene, and the next five most common are also extremely common molecular building blocks. Given these shortcomings, we do not use a scaffold split.
An alternative method is a similarity split based on clustering with fingerprints. A classic method for this is Butina Clustering (Butina, 1999). Unfortunately, it does not scale well to large datasets. We have developed an internal method that scales to large datasets. Our two step method first uses UMAP (McInnes et al., 2018) to reduce 1024 ECFP4 fingerprints (Rogers & Hahn, 2010) to length 10 float vectors. Then HDBSCAN (McInnes et al., 2017) is used to cluster compounds, in a GPU implementation from NVIDIA (cuml) this runs in a few hours on a single Tesla T4. Clusters are assigned to splits using a waterfall method, at each step assigning the largest remaining cluster to the smallest current split. The following settings were used:
UMAP(n_components=10, metric="jaccard", n_neighbors=30, min_dist=0.0, n_epochs=1000) HDBSCAN(min_cluster_size=10, min_samples=None, metric=’euclidian’, prediction_data=True, cluster_selection_method=’eom’)
Appendix F Molecular Docking Procedure
Tautomer and protonation state selection was performed using Epik Classic (Shelley et al., 2007) from Schrödinger Suite 2024-4 at pH 7.4. For each library member, a random subset of stereoisomers was enumerated using rdkit 2023.09.2 (up to 32), then 4 conformers per stereoisomer were generated using ETKDGv3 (Wang et al., 2020). Up to 8 of the lowest-energy conformers under the UFF force field (Rappé et al., 1992) (max 1 per stereoisomer) were retained, for compounds with UFF parameters for all atoms (otherwise, up to 8 random conformers). Docking was performed using the Vina scoring function (Trott & Olson, 2010) in Uni-Dock 1.1.2 (Yu et al., 2023) with exhaustiveness 128 and max_step 20, saving the top three poses per receptor. All of these 234k ligand states were docked against all six receptor models, using a fixed 20 Å docking box centered on the orthosteric pocket.
One receptor model represents the major conformation of DDR1 in available experimental structures, the DFG-out, C-helix-in PDB: 6FEX (Richter et al., 2018). Five receptor models were chosen to represent most experimental structures of MAPK14, with a variety of activation loop and P-loop conformations, PDB: 3KQ7, 3S3I, 5WJJ, 5XYY, and 6SFI (Cheng et al., 2009; Aiguadé et al., 2012; Kaieda et al., 2018; Wang et al., 2017; Röhm et al., 2019). All receptors were prepared using the Protein Preparation Wizard from Schrödinger Suite 2023-4, capping termini with neutral ends, and aligned into a common frame.
Docked poses are provided in SDF format with “molecule_hash” and “receptor” properties. Example poses are shown in Figure LABEL:fig:poses, where the hinge binding motif that is characteristic for kinases can be observed.